# MixAT: Combining Continuous and Discrete Adversarial Training for LLMs

### v1

@@ Abstract
Despite recent efforts in Large Language Models (LLMs) safety and alignment, current adversarial attacks on frontier LLMs are still able to force harmful generations consistently. Although adversarial training has been widely studied and shown to significantly improve the robustness of traditional machine learning models, its strengths and weaknesses in the context of LLMs are less understood. Specifically, while existing discrete adversarial attacks are effective at producing harmful content, training LLMs with concrete adversarial prompts is often computationally expensive, leading to reliance on continuous relaxations. As these relaxations do not correspond to discrete input tokens, such latent training methods often leave models vulnerable to a diverse set of discrete attacks. In this work, we aim to bridge this gap by introducing MixAT, a novel method that combines stronger discrete and faster continuous attacks during training. We rigorously evaluate MixAT across a wide spectrum of state-of-the-art attacks, proposing the At Least One Attack Success Rate (ALO-ASR) metric to capture the worst-case vulnerability of models. We show MixAT achieves substantially better robustness (ALO-ASR $<20\%$) compared to prior defenses (ALO-ASR $>50\%$), while maintaining a runtime comparable to methods based on continuous relaxations. We further analyze MixAT in realistic deployment settings, exploring how chat templates, quantization, low-rank adapters, and temperature affect both adversarial training and evaluation, revealing additional blind spots in current methodologies. Our results demonstrate that MixAT’s discrete-continuous defense offers a principled and superior robustness-accuracy tradeoff with minimal computational overhead, highlighting its promise for building safer LLMs. We provide our code and models at https://github.com/insait-institute/MixAT.
@@ 1 Introduction
Adversarial Training for LLMs. Inspired by successes on traditional models, adversarial training (AT) has been adapted to LLMs [5, 6, 7, 8, 9]. However, these approaches often face limitations. On the one hand, training with concrete adversarial prompts is computationally expensive, making it impractical for large models [4], while, recent cheaper methods based on continuous input embeddings [7] or latent representations [8, 9], remain vulnerable to stronger discrete attacks (Section 4).
This Work: Combining Continuous and Discrete Adversarial Training. We propose Mixed Adversarial Training (MixAT), a novel approach for AT of LLMs that combines the efficiency of continuous attacks with the resilience of discrete attack training to achieve a state-of-the-art robustness-utility trade-off. Concretely, MixAT uses continuous perturbations, like CAT [7], but applies them on top of discrete adversarial inputs, resulting in attacks that better cover the adversarial embedding space, as demonstrated in Fig. 1(a). Our extensive evaluation on diverse suits of attacks using our At Least One Attack Success Rate (ALO-ASR) metric reveals that, unlike prior defenses that are highly vulnerable (ALO-ASR > 50%), MixAT generalizes to new adversarial attacks not seen during training, establishing a new strong baseline (ALO-ASR < 20%) for robust LLMs.
@@ 2.1 Adversarial Attacks in Machine Learning
$\hat{\bm{x}}=\operatorname*{arg\,max}_{{\bm{x}}^{\prime}\in\mathcal{N}({\bm{x}})}~\log P_{\theta}(y^{*}|{\bm{x}}^{\prime}),$ (1)
@@ 2.2 Adversarial Attacks on LLMs
We consider the common scenario of an adversary targeting an auto-regressive LLM engaged in a sequence prediction task. The output of an LLM, defined by parameters ${\bm{\theta}}$, and given an input sequence ${\bm{x}}_{1:n}\in\mathcal{V}^{n}$ over vocabulary $\mathcal{V}$, is a response ${\bm{y}}_{1:m}\in\mathcal{V}^{m}$, generated by maximizing the likelihood $P_{\theta}({\bm{y}}_{i}|{\bm{x}}_{1:n},{\bm{y}}_{1:i-1})$. The adversary’s objective is to generate a malicious prompt $\hat{{\bm{x}}}$ such that the response $\hat{{\bm{y}}}$ violates predefined constraints $C$, we will elaborate on in Section 4.
@@
This perturbation $\hat{{\bm{\delta}}}$ is typically computed using iterative methods like projected gradient descent [16]. Notably, the perturbed embeddings $\hat{{\bm{e}}}_{i}={\bm{e}}_{{{\bm{x}}}_{i}}+{\bm{\delta}}_{i}$ does not generally correspond to any tokens in $\mathcal{V}^{n}$. In abuse of notation let ${\bm{x}}+{\bm{\delta}}$ refer to the perturbed embeddings, unifiying Eqs. 2 and 3:
$\hat{\bm{x}}=\operatorname*{arg\,max}_{{\bm{x}}^{\prime}\in\mathcal{B}^{p}({\bm{x}},\epsilon)}\log P_{\theta}(\hat{\bm{y}}|{\bm{x}}^{\prime})\vspace{-2mm}$ (5)
@@ 2.4 Adversarial Training for LLMs
Discrete Adversarial Training involves training the LLMs to refuse adversarial prompts generated before or during training ($\mathcal{R}(\mathcal{X})$ in Fig. 1(a)). For this, static methods, such as Rainbow Teaming [17], generate diverse adversarial examples at the start of training. In contrast, dynamic methods like R2D2 [5] and SAP (Sequential Adversarial Prompting) [18] generate adversarial samples iteratively throughout training, improving adaptability but increasing computational cost. For example, R2D2 finetunes LLMs on adversarial GCG suffixes, requiring over 100 GPU-hours for a 7B model. Adversarial Tuning [6] reduces this by precomputing candidate suffixes and only refining a subset.
Continuous Adversarial Training To mitigate the computational overhead of discrete approaches, Continuous Adversarial Training (CAT) [7] generates adversarial examples in the input embedding space ($\mathcal{X}+{\bm{\delta}}$ in Fig. 1(a)). While this method is computationally efficient and inspired by defenses in vision and audio models, most perturbed embeddings do not correspond to valid text sequences. Latent Adversarial Training (LAT) [8, 9] extends adversarial perturbations to the model’s hidden layer states. Both approaches achieve robustness against certain attacks at lower computational costs than discrete methods, though their efficacy against diverse discrete attacks remains limited (Table 1).
Prompt-Based Defenses Another complementary approach is to train a fixed prompt prefix via gradient-based optimization to steer the model away from adversarial completions [19, 20]. These methods are lightweight, model-agnostic, and effective against direct malicious requests, but they leave the model vulnerable against more sophisticated attacks (Table 1).
@@ 3.1 Mixed Adversarial Training (MixAT)
Building on Section 2 and as shown in Fig. 2, the core of MixAT lies in unifying continuous and discrete adversarial training by extending the perturbation set to include both continuous and discrete components. Following prior works, we train our models using the same principles and loss functions as described in Section 2.4. The adversarial perturbation space in MixAT is defined as:
$\mathcal{N}_{\textsc{MixAT}}({\bm{x}})=\underbrace{\mathcal{R}({\bm{x}})}_{\text{(1)}}+\underbrace{\mathcal{B}^{2}(0,\epsilon)}_{\text{(2)}}=\left\{\hat{{\bm{x}}}\mid\hat{{\bm{x}}}={\bm{x}}^{\prime}+{\bm{\delta}},{\bm{x}}^{\prime}\in\mathcal{R}({\bm{x}}),||{\bm{\delta}}||_{2}\leq\epsilon\right\},$ (8)
as represented by $\mathcal{R}(\mathcal{X})+{\bm{\delta}}$ in Fig. 1(a). Intuitively MixAT aims to center any continuous adversarial perturbation around a discrete adversarial example instead of an original benign data point.
Discrete To cover a broad range of adversarial examples in the discrete part of Eq. 8, we generate discrete adversarial seed points $\mathcal{R}({\bm{x}})$ using the adversarial training variant of PAP [2] (PAP-AT):
@@
where $\mathcal{N}_{\text{PAP}}({\bm{x}})$ represents the space of paraphrased adversarial texts around the original sample ${\bm{x}}$. While our Mixed Adversarial Training procedure is compatible with using any adversarial attack as discrete seeds, we chose to train with PAP samples since they are cheap to generate, yet very strong and diverse. In Section 4.3 we analyse a MixAT variant which also includes GCG samples. While it results in marginally better robustness, we find the 8x higher computational cost too large in practice.
Continuous For the continuous part of Eq. 8, we build upon CAT [7], optimizing model robustness in the embedding space by defining the perturbation set as an $L^{2}$ ball around the embedding of ${\bm{x}}$:
@@ 3.2 Empirical Motivation for MixAT
MixAT is designed to address the limitations of existing adversarial training methods by combining discrete and continuous attacks. Compared to DualAT and purely discrete of continuous attacks, MixAT explores a wider region of the adversarial space, leading to improved robustness against a diverse set of attacks, as visualized in Fig. 1(a). To empirically validate this, in Section B.10 we conduct a qualitative analysis of different attack prompts and their ability to induce harmful behaviors in LLMs. Moreover, in Fig. 6 we quantify the cosine similarities between different prompts in the discrete and continuous perturbation spaces. We observe that the combination of PAP and continuous attacks creates a prompt that is the least similar to the original malicious request, while also being overall closer to the GCG samples. This suggests that training using MixAT attacks can improve robustness against a wide range of attacks, including those that are not directly included in the MixAT training set. We demonstrate this empirically in Section 4, showing MixAT leads to significant gains in robustness both against direct attacks and diverse adversarial benchmarks.
@@ 4.1 Experimental Setup
We chose to evaluate MixAT on four different open-source models of varying sizes and capabilities: Zephyr-7B [24], Llama3-8B [25], Qwen2.5-14B and Qwen2.5-32B [26]. Most of our experiments, ablations, and design choices were made on Zephyr-7B, while extended evaluations on other models highlight the generalizability of our method. Unless stated otherwise, we follow the design and hyperparameter choices of prior work [7]. The default PAP sample ratio is $\alpha=0.5$, with paraphrases drawn randomly from all 40 strategies [2]. For more details on the training process see Appendix C.
@@
Model Utility Scores [%] $\uparrow$ Attack Success Rate [%] $\downarrow$
@@
Zephyr-7B No Defense (HF) 81.0 55.2 56.2 100.0 61.4 98.8 85.0 87.5 85.0 97.5 90.0 85.0 100.0 100.0
R2D2 [5] (HF) 80.1 52.9 56.1 30.0 50.2 33.6 7.5 65.0 15.0 7.5 7.5 0.0 45.0 77.5
CAT [7] (HF) 78.2 51.1 54.8 97.5 55.5 50.8 2.5 40.0 42.5 42.5 2.5 5.0 5.0 70.0
CAT [7] (R) 78.2 50.5 54.5 95.0 53.7 50.0 0.0 25.0 27.5 55.0 0.0 12.5 0.0 67.5
LAT KL [9] (R) 50.3 34.5 55.4 95.0 62.1 93.2 10.0 62.5 85.0 85.0 37.5 45.0 80.0 97.5
LAT SFT [9] (R) 31.7 23.2 22.9 45.0 54.4 38.4 5.0 30.0 30.0 27.5 2.5 20.0 15.0 52.5
PAP-AT 82.3 54.2 56.4 97.5 54.1 94.0 17.5 2.5 5.0 15.0 2.5 55.0 57.5 77.5
DualAT 81.8 54.4 56.1 85.0 54.2 47.2 2.5 2.5 10.0 15.0 0.0 10.0 2.5 22.5
MixAT 81.4 54.0 55.8 97.5 54.3 74.0 0.0 0.0 0.0 0.0 0.0 12.5 5.0 15.0
MixAT + GCG 81.6 54.5 55.9 92.5 54.6 56.4 2.5 0.0 2.5 5.0 0.0 2.5 2.5 7.5
Llama3-8B No Defense (HF) 79.1 49.1 60.8 100.0 73.3 98.0 25.0 45.0 47.5 67.5 22.5 47.5 82.5 90.0
CAT [7] (R) 79.7 50.9 58.0 65.0 65.7 48.4 0.0 30.0 47.5 70.0 0.0 7.5 5.0 82.5
LAT KL [9] (HF) 73.1 42.7 58.3 100.0 67.3 63.6 2.5 22.5 10.0 20.0 0.0 0.0 25.0 40.0
LAT KL [9] (R) 57.9 33.5 55.9 97.5 72.2 84.4 2.5 30.0 20.0 37.5 0.0 17.5 52.5 67.5
RPO [20] (GH) 71.8 42.2 54.6 100.0 58.6 97.6 17.5 35.0 15.0 35.0 5.0 60.0 100.0 100.0
PAT [19] (GH) 72.8 40.0 56.5 97.5 64.5 72.0 5.0 42.5 20.0 30.0 0.0 27.5 60.0 72.5
PAT [19] (R) 76.6 44.5 57.4 87.5 66.1 79.2 7.5 37.5 25.0 22.5 7.5 30.0 72.5 82.5
PAP-AT 81.1 51.9 60.2 100.0 58.6 84.4 22.5 2.5 15.0 22.5 10.0 52.5 40.0 70.0
DualAT 80.7 50.6 59.9 67.5 57.1 32.8 0.0 10.0 7.5 25.0 0.0 20.0 0.0 37.5
MixAT 80.4 50.1 59.1 85.0 55.6 40.0 0.0 0.0 2.5 2.5 0.0 22.5 0.0 25.0
MixAT + GCG 80.1 48.7 58.5 92.5 55.2 47.6 0.0 0.0 5.0 7.5 0.0 2.5 0.0 15.0
Qwen2.5-14B No Defense (HF) 83.8 57.8 77.7 100.0 85.3 99.2 15.0 57.5 75.0 82.5 37.5 70.0 100.0 100.0
CAT [7] (R) 84.9 59.7 76.5 92.5 71.8 52.4 2.5 30.0 62.5 72.5 0.0 5.0 2.5 92.5
LAT KL [9] (R) 81.8 54.4 72.7 95.0 82.6 78.8 2.5 30.0 30.0 42.5 0.0 27.5 2.5 75.0
PAP-AT 84.8 58.7 77.1 95.0 71.8 71.6 5.0 10.0 27.5 50.0 0.0 67.5 2.5 85.0
MixAT 86.2 60.8 75.6 90.0 62.3 40.4 0.0 0.0 5.0 7.5 0.0 5.0 0.0 15.0
MixAT + GCG 84.2 59.5 75.8 87.5 66.5 46.8 0.0 0.0 2.5 5.0 0.0 2.5 0.0 5.0
Qwen-32B No Defense (HF) 82.5 57.8 81.1 100.0 85.4 98.8 10.0 60.0 87.5 97.5 17.5 70.0 100.0 100.0
CAT [7] (R) 83.1 58.8 79.7 92.5 72.4 42.0 0.0 27.5 40.0 65.0 0.0 12.5 0.0 82.5
PAP-AT 84.0 58.9 80.8 100.0 71.6 92.0 17.5 35.0 80.0 92.5 55.0 80.0 100.0 100.0
MixAT 83.9 59.4 80.7 90.0 70.3 47.2 0.0 0.0 0.0 2.5 0.0 7.5 0.0 7.5
@@
Evaluation Metrics Our goal is to train models that are both robust against adversarial attacks and maintain high utility. To assess robustness, we use the Attack Success Rate (ASR), which quantifies the percentage of adversarial samples that successfully induce harmful model responses to malicious inputs. We evaluate against a variety of adversarial methods, including PAP [2], TAP [10], PAIR [3], AutoDAN [11], GCG [4], and HumanJailbreaks [27]. Additionally, we test the model’s resistance to direct malicious requests not tied to specific attack methods. Following prior work [5, 7, 9, 6], we use the HarmBench dataset [5]. Since small models (7-8B parameters) often struggle with reproducing copyright content, we restrict our evaluation to the first 40 non-copyright-related samples in the HarmBench test set (details in Section B.4). Since different attack strategies often succeed on different samples, we also report the "At Least One" Attack Success Rate (ALO-ASR), reflecting the success rate of a meta-adversary using all attacks. ALO-ASR serves as a proxy for universal robustness.

### v2

@@ Abstract
Despite recent efforts in Large Language Model (LLM) safety and alignment, current adversarial attacks on frontier LLMs can still consistently force harmful generations. Although adversarial training has been widely studied and shown to significantly improve the robustness of traditional machine learning models, its strengths and weaknesses in the context of LLMs are less understood. Specifically, while existing discrete adversarial attacks are effective at producing harmful content, training LLMs with concrete adversarial prompts is often computationally expensive, leading to reliance on continuous relaxations. At the same time, despite their effectiveness and generalization capabilities, training with continuous perturbations does not always capture the full spectrum of vulnerabilities exploited by discrete attacks. In this work, we aim to bridge this gap by introducing MixAT, a novel method that combines stronger discrete and faster continuous attacks during training. We rigorously evaluate MixAT across a wide spectrum of state-of-the-art attacks, proposing the At Least One Attack Success Rate (ALO-ASR) metric to capture the worst-case vulnerability of models. We show MixAT achieves substantially better robustness (ALO-ASR $20\%$) compared to prior defenses (ALO-ASR $50\%$), while maintaining a runtime comparable to methods based on continuous relaxations. We further analyze MixAT in realistic deployment settings, exploring how chat templates, quantization, low-rank adapters, and temperature affect both adversarial training and evaluation, revealing additional blind spots in current methodologies. Our results demonstrate that MixAT’s discrete-continuous defense offers a principled and superior robustness-accuracy tradeoff with minimal computational overhead, highlighting its promise for building safer LLMs. We provide our code and models at https://github.com/insait-institute/MixAT.
@@ 1 Introduction
Adversarial Training for LLMs. Inspired by successes on traditional models, adversarial training (AT) has been adapted to LLMs [5, 6, 7, 8, 9]. However, these approaches often face limitations. On the one hand, training with concrete adversarial prompts is computationally expensive, making it impractical for large models [4], while, recent cheaper methods based on continuous input embeddings [7] or latent representations [8, 9], remain vulnerable to stronger discrete attacks (Section˜4).
This Work: Combining Continuous and Discrete Adversarial Training. We propose Mixed Adversarial Training (MixAT), a novel approach for AT of LLMs that combines the efficiency of continuous attacks with the resilience of discrete attack training to achieve a state-of-the-art robustness-utility trade-off. Concretely, MixAT uses continuous perturbations, like CAT [7], but applies them on top of discrete adversarial inputs, resulting in attacks that better cover the adversarial embedding space, as demonstrated in Fig.˜1(a). Our extensive evaluation on diverse suits of attacks using our At Least One Attack Success Rate (ALO-ASR) metric reveals that, unlike prior defenses that are highly vulnerable (ALO-ASR > 50%), MixAT generalizes to new adversarial attacks not seen during training, establishing a new strong baseline (ALO-ASR < 20%) for robust LLMs.
@@ 2.1 Adversarial Attacks in Machine Learning
$\hat{\bm{x}}=\operatorname*{arg\,max}_{{\bm{x}}^{\prime}\in\mathcal{N}({\bm{x}})} \log P_{\theta}(y^{*}|{\bm{x}}^{\prime}),$ (1)
@@ 2.2 Adversarial Attacks on LLMs
We consider the common scenario of an adversary targeting an auto-regressive LLM engaged in a sequence prediction task. The output of an LLM, defined by parameters $\bm{\theta}$}, and given an input sequence ${\bm{x}}_{1:n}\in\mathcal{V}^{n}$ over vocabulary $\mathcal{V}$, is a response ${\bm{y}}_{1:m}\in\mathcal{V}^{m}$, generated by maximizing the likelihood $P_{\theta}({\bm{y}}_{i}|{\bm{x}}_{1:n},{\bm{y}}_{1:i-1})$. The adversary’s objective is to generate a malicious prompt $\hat{{\bm{x}}}$ such that the response $\hat{{\bm{y}}}$ violates predefined constraints $C$, we will elaborate on in Section˜4.
@@
This perturbation $\hat{{\bm{\delta}}}$ is typically computed using iterative methods like projected gradient descent [16]. Notably, the perturbed embeddings $\hat{{\bm{e}}}_{i}={\bm{e}}_{{{\bm{x}}}_{i}}+{\bm{\delta}}_{i}$ does not generally correspond to any tokens in $\mathcal{V}^{n}$. In abuse of notation let ${\bm{x}}+{\bm{\delta}}$ refer to the perturbed embeddings, unifiying Eqs.˜2 and 3:
$\hat{\bm{x}}=\operatorname*{arg\,max}_{{\bm{x}}^{\prime}\in\mathcal{B}^{p}({\bm{x}},\epsilon)}\log P_{\theta}(\hat{\bm{y}}|{\bm{x}}^{\prime})\vskip-5.69054pt$ (5)
Model Tampering attacks Besides input space and latent space attacks, model tampering attacks allow modification of weights and latent activations [17]. For this work, we consider such stronger adversarial attacks out of scope, but consider them an interesting avenue for future work.
@@ 2.4 Adversarial Training for LLMs
Discrete Adversarial Training involves training the LLMs to refuse adversarial prompts generated before or during training ($\mathcal{R}(\mathcal{X})$ in Fig.˜1(a)). For this, static methods, such as Rainbow Teaming [18], generate diverse adversarial examples at the start of training. In contrast, dynamic methods like R2D2 [5] and SAP (Sequential Adversarial Prompting) [19] generate adversarial samples iteratively throughout training, improving adaptability but increasing computational cost. For example, R2D2 finetunes LLMs on adversarial GCG suffixes, requiring over 100 GPU-hours for a 7B model. Adversarial Tuning [6] reduces this by precomputing candidate suffixes and only refining a subset.
Continuous Adversarial Training To mitigate the computational overhead of discrete approaches, Continuous Adversarial Training (CAT) [7] generates adversarial examples in the input embedding space ($\mathcal{X}+{\bm{\delta}}$ in Fig.˜1(a)). While this method is computationally efficient and inspired by defenses in vision and audio models, most perturbed embeddings do not correspond to valid text sequences. Latent Adversarial Training (LAT) [8, 9] extends adversarial perturbations to the model’s hidden layer states. Both approaches achieve robustness against certain attacks at lower computational costs than discrete methods, though their efficacy against diverse discrete attacks remains limited (Table˜1).
Prompt-Based Defenses Another complementary approach is to train a fixed prompt prefix via gradient-based optimization to steer the model away from adversarial completions [20, 21]. These methods are lightweight, model-agnostic, and effective against direct malicious requests, but they leave the model vulnerable against more sophisticated attacks (Table˜1).
@@ 2.5 Other Defenses
One particularly prominent direction in model-augmented defenses is the principle of Constitutional AI, wherein humans specify a set of normative guidelines that govern a model’s behavior to mitigate harmful outputs [25, 26]. Under these guidelines, model (self-)critique can be employed to generate additional training data or serve as reinforcement learning feedback [25], promoting safer and more aligned behavior. Constitutional Classifiers [26] build on this approach by leveraging synthetic training data aligned with natural-language constitutions to monitor both model inputs and outputs for harmful content. While such classifiers provide an additional layer of defense against a broad range of attacks, we consider them orthogonal to adversarial model training.
@@ 3.1 Mixed Adversarial Training (MixAT)
Building on Section˜2 and as shown in Fig.˜2, the core of MixAT lies in unifying continuous and discrete adversarial training by extending the perturbation set to include both continuous and discrete components. Following prior works, we train our models using the same principles and loss functions as described in Section˜2.4. The adversarial perturbation space in MixAT is defined as:
[…]
as represented by $\mathcal{R}(\mathcal{X})+{\bm{\delta}}$ in Fig.˜1(a). Intuitively MixAT aims to center any continuous adversarial perturbation around a discrete adversarial example instead of an original benign data point.
Discrete To cover a broad range of adversarial examples in the discrete part of Eq.˜8, we generate discrete adversarial seed points $\mathcal{R}({\bm{x}})$ using the adversarial training variant of PAP [2] (PAP-AT):
@@
where $\mathcal{N}_{\text{PAP}}({\bm{x}})$ represents the space of paraphrased adversarial texts around the original sample $\bm{x}$}. While our Mixed Adversarial Training procedure is compatible with using any adversarial attack as discrete seeds, we chose to train with PAP samples since they are cheap to generate, yet very strong and diverse. In Section˜4.3 we analyse a MixAT variant which also includes GCG samples. While it results in marginally better robustness, we find the 8x higher computational cost too large in practice.
Continuous For the continuous part of Eq.˜8, we build upon CAT [7], optimizing model robustness in the embedding space by defining the perturbation set as an $L^{2}$ ball around the embedding of $\bm{x}$}:
@@ 3.2 Empirical Motivation for MixAT
MixAT is designed to address the limitations of existing adversarial training methods by combining discrete and continuous attacks. Compared to DualAT and purely discrete of continuous attacks, MixAT explores a wider region of the adversarial space, leading to improved robustness against a diverse set of attacks, as visualized in Fig.˜1(a). To empirically validate this, in Section˜B.14 we conduct a qualitative analysis of different attack prompts and their ability to induce harmful behaviors in LLMs. Moreover, in Fig.˜6 we quantify the cosine similarities between different prompts in the discrete and continuous perturbation spaces. We observe that the combination of PAP and continuous attacks creates a prompt that is the least similar to the original malicious request, while also being overall closer to the GCG samples. This suggests that training using MixAT attacks can improve robustness against a wide range of attacks, including those that are not directly included in the MixAT training set. We demonstrate this empirically in Section˜4, showing MixAT leads to significant gains in robustness both against direct attacks and diverse adversarial benchmarks.
@@ 4.1 Experimental Setup
We chose to evaluate MixAT on four different open-source models of varying sizes and capabilities: Zephyr-7B [27], Llama3-8B [28], Qwen2.5-14B and Qwen2.5-32B [29]. Most of our experiments, ablations, and design choices were made on Zephyr-7B, while extended evaluations on other models highlight the generalizability of our method. Unless stated otherwise, we follow the design and hyperparameter choices of prior work [7]. The default PAP sample ratio is $\alpha=0.5$, with paraphrases drawn randomly from all 40 strategies [2]. For more details on the training process see Appendix˜C.
@@
Model
Utility Scores [%] $\uparrow$ Attack Success Rate [%] $\downarrow$
@@
Zephyr-7B
No Defense (HF) 81.0 55.2 56.2 100.0 60.3 98.8 85.0 87.5 85.0 97.5 90.0 85.0 100.0 100.0
R2D2 [5] (HF) 80.1 52.9 56.1 30.0 42.2 33.6 7.5 65.0 15.0 7.5 7.5 0.0 45.0 77.5
CAT [7] (HF) 78.2 51.1 54.8 97.5 52.8 50.8 2.5 40.0 42.5 42.5 2.5 5.0 5.0 70.0
CAT [7] (R) 78.2 50.5 54.5 95.0 52.3 50.0 0.0 25.0 27.5 55.0 0.0 12.5 0.0 67.5
LAT KL [9] (R) 50.3 34.5 55.4 95.0 60.9 93.2 10.0 62.5 85.0 85.0 37.5 45.0 80.0 97.5
LAT SFT [9] (R) 31.7 23.2 22.9 45.0 32.6 38.4 5.0 30.0 30.0 27.5 2.5 20.0 15.0 52.5
PAP-AT 82.3 54.2 56.4 97.5 52.6 94.0 17.5 2.5 5.0 15.0 2.5 55.0 57.5 77.5
DualAT 81.8 54.4 56.1 85.0 51.1 47.2 2.5 2.5 10.0 15.0 0.0 10.0 2.5 22.5
MixAT 81.4 54.0 55.8 97.5 52.2 74.0 0.0 0.0 0.0 0.0 0.0 12.5 5.0 15.0
MixAT + GCG 81.6 54.5 55.9 92.5 51.1 56.4 2.5 0.0 2.5 5.0 0.0 2.5 2.5 7.5
Llama3-8B
No Defense (HF) 79.1 49.1 60.8 100.0 68.6 98.0 25.0 45.0 47.5 67.5 22.5 47.5 82.5 90.0
CAT [7] (R) 79.7 50.9 58.0 65.0 61.6 48.4 0.0 30.0 47.5 70.0 0.0 7.5 5.0 82.5
LAT KL [9] (HF) 73.1 42.7 58.3 100.0 67.1 63.6 2.5 22.5 10.0 20.0 0.0 0.0 25.0 40.0
LAT KL [9] (R) 57.9 33.5 55.9 97.5 69.4 84.4 2.5 30.0 20.0 37.5 0.0 17.5 52.5 67.5
RPO [21] (GH) 71.8 42.2 54.6 100.0 45.1 97.6 17.5 35.0 15.0 35.0 5.0 60.0 100.0 100.0
PAT [20] (GH) 72.8 40.0 56.5 97.5 69.2 72.0 5.0 42.5 20.0 30.0 0.0 27.5 60.0 72.5
PAT [20] (R) 76.6 44.5 57.4 87.5 67.9 79.2 7.5 37.5 25.0 22.5 7.5 30.0 72.5 82.5
PAP-AT 81.1 51.9 60.2 100.0 57.5 84.4 22.5 2.5 15.0 22.5 10.0 52.5 40.0 70.0
DualAT 80.7 50.6 59.9 67.5 54.9 32.8 0.0 10.0 7.5 25.0 0.0 20.0 0.0 37.5
MixAT 80.4 50.1 59.1 85.0 53.1 40.0 0.0 0.0 2.5 2.5 0.0 22.5 0.0 25.0
MixAT + GCG 80.1 48.7 58.5 92.5 50.7 47.6 0.0 0.0 5.0 7.5 0.0 2.5 0.0 15.0
Qwen2.5-14B
No Defense (HF) 83.8 57.8 77.7 100.0 83.0 99.2 15.0 57.5 75.0 82.5 37.5 70.0 100.0 100.0
CAT [7] (R) 84.9 59.7 76.5 92.5 70.8 52.4 2.5 30.0 62.5 72.5 0.0 5.0 2.5 92.5
LAT KL [9] (R) 81.8 54.4 72.7 95.0 82.4 78.8 2.5 30.0 30.0 42.5 0.0 27.5 2.5 75.0
PAP-AT 84.8 58.7 77.1 95.0 69.1 71.6 5.0 10.0 27.5 50.0 0.0 67.5 2.5 85.0
MixAT 86.2 60.8 75.6 90.0 64.9 40.4 0.0 0.0 5.0 7.5 0.0 5.0 0.0 15.0
MixAT + GCG 84.2 59.5 75.8 87.5 64.6 46.8 0.0 0.0 2.5 5.0 0.0 2.5 0.0 5.0
Qwen2.5-32B
No Defense (HF) 82.5 57.8 81.1 100.0 85.2 98.8 10.0 60.0 87.5 97.5 17.5 70.0 100.0 100.0
CAT [7] (R) 83.1 58.8 79.7 92.5 72.9 42.0 0.0 27.5 40.0 65.0 0.0 12.5 0.0 82.5
PAP-AT 84.0 58.9 80.8 100.0 70.8 92.0 17.5 35.0 80.0 92.5 55.0 80.0 100.0 100.0
MixAT 83.9 59.4 80.7 90.0 66.9 47.2 0.0 0.0 0.0 2.5 0.0 7.5 0.0 7.5
@@
Evaluation Metrics Our goal is to train models that are both robust against adversarial attacks and maintain high utility. To assess robustness, we use the Attack Success Rate (ASR), which quantifies the percentage of adversarial samples that successfully induce harmful model responses to malicious inputs. We evaluate against a variety of adversarial methods, including PAP [2], TAP [10], PAIR [3], AutoDAN [11], GCG [4], and HumanJailbreaks [30]. Additionally, we test the model’s resistance to direct malicious requests not tied to specific attack methods. Following prior work [5, 7, 9, 6], we use the HarmBench dataset [5]. Since small models (7-8B parameters) often struggle with reproducing copyright content, we restrict our evaluation to the first 40 non-copyright-related samples in the HarmBench test set (details in Section˜B.4). Since different attack strategies often succeed on different samples, we also report the "At Least One" Attack Success Rate (ALO-ASR), reflecting the success rate of a meta-adversary using all attacks. ALO-ASR serves as a proxy for universal robustness. We obtain a score of $18.75\%$ on a scale from 0 (worst) to 100 (best).